# Effective Online Exam Proctoring by Combining Lightweight Face Detection and Deep Recognition

### v2

@@ Abstract
Online exams, conducted via video conferencing platforms such as Zoom, have become popular in educational institutions since COVID-19. While convenient, ensuring the integrity and security of online exams remains challenging, as traditional invigilation struggles to effectively monitor multiple student video feeds in real time. In this paper, we present iExam, an effective online exam proctoring and analysis system that combines lightweight face detection and deep recognition. iExam employs real-time face detection to assist invigilators in continuously monitoring student presence, and leverages deep face recognition for post-exam video analysis to identify abnormal behaviors—including face disappearance, face rotation, and identity substitution. To realize this system, we address three core challenges: (i) designing a lightweight approach to efficiently capture and analyze exam video streams in real time; (ii) developing an enhanced OCR method to automatically extract student identities from dynamically positioned Zoom name tags, enabling reliable ground truth labeling without manual intervention; and (iii) optimizing the training and inference pipeline to significantly reduce resource and time requirements on ordinary teacher devices. Extensive experiments demonstrate that iExam achieves 90.4% accuracy for real-time face detection and 98.4% accuracy for post-exam face recognition, while maintaining low overhead. These results show that iExam can substantially enhance the automation and reliability of online exam proctoring in practice.
@@ I Introduction
The rapid adoption of online and hybrid education, especially in the wake of the COVID-19 pandemic, has fundamentally changed the landscape of teaching and assessment [chakraborty2021opinion, bozkurt2022impact, pokhrel2021literature]. Synchronous online exams delivered over video conferencing platforms such as Zoom and Microsoft Teams have become commonplace in schools and universities worldwide [linden2021zoom, microsoft_teams_exam]. While online exams offer great flexibility and accessibility, ensuring their integrity and fairness remains a critical challenge [butler2020systematic, standard_online_exam_cheats, swissinfo_online_exam_concerns, holden2021academic, langenfeld2020internet].
A central difficulty lies in the effective monitoring—or invigilation—of remote exams. Unlike traditional classroom settings, where instructors can physically oversee all students’ behavior, online environments require proctors to supervise dozens of simultaneous video streams, typically presented in a crowded grid of small windows [lee2022online]. Manual monitoring is inherently limited by human attention and is prone to oversight, bias and fatigue, especially during large-scale or extended assessments [belzak2024measuring].
Existing approaches to online proctoring fall broadly into two categories. Commercial automated proctoring systems leverage computer vision and AI to analyze student behavior, but they often depend on intrusive data collection, centralized cloud processing, or require complex installation and configuration [proctoru_official_website, balash2021examining, balash2023educators]. These factors raise privacy concerns and limit adoption in privacy-sensitive or resource-constrained environments. In contrast, manual video monitoring remains widely used but is inherently unscalable and susceptible to subjective errors [pongchaikul2024remote].
To address these limitations, we present iExam, a novel and practical online exam proctoring system that combines lightweight face detection for real-time monitoring with deep learning-based face recognition for post-exam analysis. The core insight behind iExam is to decouple instant invigilation from in-depth integrity checks: lightweight detection runs efficiently in the browser for instant feedback, while deep recognition is deferred to offline processing for maximum accuracy and flexibility.
During the exam, iExam employs modern Web APIs to capture the Zoom gallery view and perform face detection entirely on the client side [getdisplaymedia_api, opencvjs], thereby preserving privacy and enabling scalable real-time performance without transmitting sensitive images to external servers. After the exam, iExam’s post-processing module leverages state-of-the-art Deep Convolutional Neural Networks (DCNNs) [krizhevsky2012imagenet, he2016deep] to recognize faces and flag anomalies, such as face disappearance, rotation, or identity substitution, from recorded video streams. To automate the collection of ground truth labels for training, we further introduce an enhanced Optical Character Recognition (OCR) technique that reliably extracts student names from dynamic on-screen labels [tesseract_github, yaqub2023proctoring].
To ensure that iExam remains accessible and practical for real-world deployment, for example on standard teacher laptops, which often involves limited computational resources, we incorporate a series of efficiency-oriented optimizations that significantly reduce training and computational overhead, without sacrificing accuracy. Key techniques include sparse OCR labeling, which minimizes redundant text extraction operations; multi-threaded processing, enabling smooth performance even on modest hardware; and automated model selection, which dynamically balances speed and performance based on available resources. These optimizations allow robust face detection algorithms and deep convolutional neural networks to be executed on lightweight hardware, eliminating the need for dedicated GPU servers or costly cloud infrastructure. By keeping both detection and recognition pipelines locally deployable, iExam offers a privacy-preserving and resource-efficient solution for online exam proctoring.
We validate iExam through a series of comprehensive experiments on real-world exam datasets collected from university courses. Results show that our system consistently achieves high accuracy in both real-time face detection (over 90%) and post-exam recognition (about 98%), with acceptable runtime performance and scalability. Moreover, instructors benefit from automated visual reports and anomaly summaries that not only reduce manual monitoring workload but also provide auditable exam integrity checks, improving trust and transparency in online assessments. These results highlight that iExam is a deployable and effective solution bridging the gap between lightweight real-time proctoring and rigorous post-exam analysis.
Contributions. In sum, our contributions are summarized as follows:
@@
  • We design and implement iExam, a two-stage online exam proctoring system that decouples lightweight, privacy-preserving real-time face detection from deep learning-based post-exam face recognition and anomaly analysis.
@@
  • We develop an automated ground truth labeling pipeline using enhanced OCR techniques, enabling reliable extraction of student identities from dynamic Zoom window labels without manual annotation.
  • We introduce a set of resource-efficient training and inference optimizations, including sparse OCR labeling, batch and multi-threaded processing, and model selection strategies, allowing our system to run effectively on standard teacher hardware.
@@
  • We conduct comprehensive experiments with real-world online exam videos, demonstrating that iExam achieves high accuracy (over 90% in real-time detection and 98% in post-exam recognition) and delivers interpretable, actionable reports to instructors for scalable and trustworthy online exam supervision.
II Background
In this section, we introduce the typical online exam scenario using Zoom as the platform, and review mainstream face detection and recognition algorithms.
II-A Online Exams in the “Zoom” Setting
Zoom is one of the most widely adopted platforms for online education and remote examinations [linden2021zoom], providing real-time video conferencing and recording functionalities. In its “gallery view” mode, Zoom displays up to 25 participants per screen (5 rows $\times$ 5 columns) by default, and up to 49 participants (7 $\times$ 7) with modified settings. When the number of exam-takers exceeds the single-screen capacity, additional assistants are required to record multiple screens in parallel. Each session can be recorded and saved as an “mp4” video file, preserving all participants’ facial information, which serves as the input for our iExam system.
A unique feature of Zoom is the personalized label—each participant’s name is shown in the lower left corner of their video window (see Fig. 1), which provides a potential anchor for automated identity extraction. While Zoom enables invigilators to monitor students’ surroundings, it lacks built-in algorithms to detect face disappearance, abnormal behavior, or potential identity substitution, necessitating the development of automated proctoring solutions.
Figure 1: Zoom gallery view with participant video feeds and name labels (masked for anonymity), serving as the input for automated face detection and identity recognition in iExam.
II-B Mainstream Face Classification Approaches
Over the years, researchers have proposed a variety of algorithms for face classification, which can be broadly categorized into feature-based and image-based methods [kumar2019face].
Feature-based approaches exploit properties such as gray-scale values, color, or edges for face detection. In the image-based category, linear subspace techniques, such as Eigenface [turk1991eigenfaces, turk1991face], Principal Component Analysis (PCA) [sirovich1987low, ramadhani2017human], and Linear Discriminant Analysis (LDA) [chan2008face], have shown promise for face recognition by projecting facial images into lower-dimensional subspaces.
However, recent advances in deep learning, especially convolutional neural networks (CNNs) [howard2017mobilenets, huang2017densely, SqueezeNet, szegedy2015going], have significantly improved both the accuracy and robustness of face detection and recognition. Deep Convolutional Neural Networks (DCNNs) [lecun1998gradient, krizhevsky2012imagenet, he2016deep] represent a more advanced class of models that extend traditional CNN architectures to achieve higher accuracy and robustness. Some models [deng2020retinaface, zhang2016joint, wang2023efficientface], pre-trained on large-scale face datasets, achieve lower detection error rates and faster inference compared to traditional algorithms, making them especially suitable for automated online exam proctoring.
III System Overview
In this section, we present a high-level overview of iExam, highlighting its core objectives, architectural rationale, and system workflow. Fig. 2 illustrates the overall architecture.
@@ III System Overview
The key design insight behind iExam is to combine lightweight face detection for real-time monitoring with deep face recognition for accurate, retrospective analysis, achieving both efficiency and robustness for large-scale online exam proctoring. This two-phase architecture is motivated by the distinct requirements at different stages of an online exam:
  • During the exam, real-time proctoring requires rapid feedback on student presence and attention. Here, a lightweight face detection module runs efficiently in the browser, enabling invigilators to promptly identify missing or inattentive students, while minimizing system latency and resource consumption.
  • After the exam, post-exam analysis demands greater accuracy and richer behavioral assessment. Full exam recordings are processed offline using deep learning-based face recognition, which can robustly verify student identities and detect sophisticated anomalies such as face disappearance, rotation, or substitution.
Based on this insight, iExam is composed of two tightly-coupled components:
1) Real-Time Proctoring: iExam provides a web-based platform that captures the Zoom gallery view and applies fast face detection to each video feed in real time. Its primary purpose is to deliver immediate warnings to invigilators — for example, alerting them when a student’s face is absent or turned away from the screen, so that potential issues can be addressed promptly on the spot while efficiently monitoring multiple students simultaneously.
2) Post-Exam Analysis: After the exam, iExam extracts face data from a short training video sampled from the tested video, trains a deep recognition model for each student, and then analyzes the entire recorded session. Unlike real-time proctoring, this component focuses on more in-depth detection of sophisticated anomalies, such as identity substitution. It then generates detailed summary reports, enabling retrospective review and further investigation.
This two-phase design allows iExam to support both real-time invigilation and retrospective integrity verification, providing a robust solution for scalable online exam proctoring.
IV Lightweight Real-time Video Monitoring based on Face Detection
@@
A key advantage of iExam is its privacy-preserving, fully client-side design. All face detection is performed locally within the browser—no video frames or facial features are ever transmitted externally. This edge-based architecture eliminates privacy risks inherent to cloud processing and enables seamless, scalable deployment in real-world educational settings.
Figure 3: Web-based proctoring interface: the invigilator selects the Zoom gallery window for monitoring using the browser’s screen capture feature.
Figure 4: Real-time face detection result: detected face regions are overlaid on the mirrored Zoom feed in the browser, providing instant feedback for invigilators.
User Workflow. During an online exam, invigilators open the web-based iExam platform and select the Zoom window to monitor (Fig. 3). The system mirrors the chosen window and overlays real-time face detection results, displaying both the original and processed video streams (Fig. 4). The face detection model is loaded on demand, and all computation occurs locally, ensuring fast and seamless operation without external dependencies.
@@
V Comprehensive Post-exam Video Analysis based on Face Recognition
@@
The workflow of this component consists of three main stages: (1) face dataset construction, (2) classifier training, and (3) face monitoring and feedback generation. To initiate the process, the proctor provides a short sample video (e.g., five minutes), which serves as the training set for individualized face model building.
@@
In this task, two primary challenges must be addressed:
@@
V-A Automatic Ground Truth Collection
@@
The Limitation of Static Position Labeling. The Zoom gallery mode arranges participants in a $5\times 5$ grid, but their order is not fixed. As students join, leave, or lose connection, the positions are shuffled—newcomers are usually placed at the lower right, while others move to fill vacant spots. Thus, it is impossible to rely on a static mapping between cell position and student identity.
@@
V-B Effective and Efficient Face Training and Recognition
@@
Model Training and Selection. We evaluated a range of state-of-the-art Convolutional Neural Networks(CNN) architectures, including AlexNet [krizhevsky2012imagenet], GoogleNet [szegedy2015going], various ResNet variants (ResNet18, ResNet50, ResNet152) [he2016deep], as well as SqueezeNet [SqueezeNet], MobileNet [howard2017mobilenets], and DenseNet121 [huang2017densely], using cross-entropy loss as the training objective. To mitigate overfitting arising from the temporal coherence of consecutive video frames, we performed data augmentation such as random flipping, rotation, and cropping during training. This increases the diversity of training samples and enhances the model’s generalization ability.
@@
VI Evaluation
VI-A The Performance and Accuracy of Face Detection Algorithms in Real-time Exam Monitoring
@@
Experimental Setup. We collected data from an actual online exam session, resulting in a two-hour video featuring the Zoom gallery view, as depicted in Fig. 3 and 4 (see Sec. IV). For this experiment, we implemented two versions of iExam’s real-time monitoring system: one using the traditional Haar-cascade face detector, and the other employing a modern DCNN-based detector. Apart from the face detection module, all other components and settings, including browser environment and platform configurations, remained identical.
@@
Runtime Performance. We first measured the detection efficiency of each face detection algorithm, focusing on the number of successful detections and the corresponding processing time. As shown in Fig. 8(a), Haar-cascade achieves about twice the detection rate of DCNN on the same video stream, primarily due to its simpler, feature-based approach and lower computational requirements. However, this higher speed comes at the expense of accuracy and robustness, as the Haar-cascade method is more prone to missing faces under challenging conditions (e.g., large head rotations, partial occlusions).
In contrast, while the DCNN-based detector processes frames at a lower rate—requiring roughly twice as long to analyze each image—it consistently provides more reliable detection results, especially in cases where faces are tilted, partially covered, or presented at non-frontal angles. This robustness is attributed to the learning-based nature of DCNNs, which enables them to generalize beyond the fixed feature templates used in traditional methods.
Detection Accuracy. To quantify detection accuracy, we conducted a manual evaluation by randomly sampling 50 frames from the two-hour video and annotating the ground truth for each student’s face. Fig. 8(b) illustrates the results, showing that the DCNN detector achieves approximately 90% detection accuracy, significantly outperforming the Haar-cascade method. Notably, the DCNN approach is better able to adapt to a variety of challenging viewing conditions, such as faces appearing close to the camera, rotated, or partially obscured. In contrast, Haar-cascade often fails to detect faces that deviate from standard frontal poses or are occluded, and sometimes misidentifies background objects as faces.
@@
VI-B The Performance and Accuracy of Face Recognition Models in Post-exam Video Analysis
@@
Experimental Setup. We first partition the available face image dataset into three disjoint subsets: 70% for training, 20% for validation, and the remaining 10% for testing. All file indices are randomly shuffled prior to splitting to avoid selection bias. To further verify the robustness of our models, we also evaluate on an independent 90-minute exam video not used during training. The distribution of the training and validation samples is illustrated in Fig. 9. All training experiments were performed on a laptop equipped with an NVIDIA GTX 1650Ti GPU (4GB VRAM), simulating a practical deployment environment for educational institutions.
Figure 9: The distribution of training and validation datasets. Real names were masked and replaced with standardized labels to preserve privacy.
Training Performance. We systematically compare the training time and recognition accuracy of eight popular DCNN architectures, including AlexNet, GoogleNet, ResNet18, ResNet50, ResNet152, SqueezeNet, MobileNet, and DenseNet121. As shown in Fig. 10, GoogleNet stands out as the fastest model, requiring only 244 minutes to complete 10 epochs of training on our dataset. AlexNet and ResNet18 also demonstrate relatively fast training speeds, completing in 293 and 303 minutes, respectively. In contrast, deeper architectures such as ResNet152 and SqueezeNet require substantially more time—650 and 555 minutes, respectively—reflecting the increased computational complexity of larger models. MobileNet and DenseNet121 fall in the mid-to-upper range, with training times of 469 and 445 minutes, respectively.
@@
Recognition Accuracy. Despite the differences in training efficiency, we observe notable variation in the face recognition accuracy achieved by each model, as summarized in the right-hand section of Fig. 10. ResNet50 achieves the highest accuracy at 98.41%, establishing a clear lead over other models. ResNet152 and GoogleNet also perform well, reaching 93.02% and 88.28% accuracy, respectively. AlexNet and MobileNet deliver moderate results, with accuracies of 77.26% and 78.33%. SqueezeNet lags behind with the lowest accuracy at 69.07%, likely due to its lightweight architecture and reduced parameter count. DenseNet121 and ResNet18 achieve intermediate performance (86.17% and 83.82%, respectively).
@@
VI-C A Case Study of iExam’s Post-exam Analysis Report
To illustrate the practical utility of iExam ’s post-exam analysis module, we present a representative case study based on a five-minute exam video. The system automatically generates a set of visual and histogram-style reports, allowing instructors to review exam-taker attendance and identify abnormal behaviors at a glance.
Recognition Frequency Overview. As shown in Fig. 11, the report includes a bar chart summarizing the total number of times each student’s face was correctly recognized during the analysis window. Although the recognition times is computed at the frame level, to improve the testing efficiency, as detailed in Sec. V-B, we treat one successful recognition per student per second as a valid presence record. Since one second typically corresponds to 30 frames, this approach significantly reduces computation while still providing a reliable measure of student presence. The visualization offers a high-level summary of engagement and can help quickly identify students whose faces were rarely detected, potentially flagging technical issues or suspicious behavior.
@@ VI-C A Case Study of iExam’s Post-exam Analysis Report
Figure 11: Conclusion of Recognition Times. The number of times each student’s face was detected and identified over the test interval. To protect privacy, the real names were replaced with generic labels.

### v3

@@ Abstract
Online exams conducted via video conferencing platforms such as Zoom have become widespread, yet ensuring exam integrity remains challenging due to the difficulty of monitoring multiple video feeds in real time. We present iExam, an online exam proctoring and analysis system that combines lightweight real-time face detection with deep face recognition for post-exam analysis. iExam assists invigilators by monitoring student presence during exams and identifies abnormal behaviors, such as face disappearance, face rotation, and identity substitution, from recorded videos. The system addresses three key challenges: (i) efficient real-time video capture and analysis, (ii) automated student identity labeling using enhanced OCR on dynamic Zoom name tags, and (iii) resource-efficient training and inference on standard teacher devices. Extensive experiments show that iExam achieves 90.4% accuracy in real-time face detection and 98.4% accuracy in post-exam recognition with low overhead, demonstrating its practicality and effectiveness for online exam proctoring.
@@ I Introduction
The rapid adoption of online and hybrid education, accelerated by the COVID-19 pandemic, has fundamentally reshaped teaching and assessment practice worldwide. Synchronous online examinations conducted via video conferencing platforms such as Zoom and Microsoft Teams are now commonplace in schools and universities. While these formates offer flexibility and accessibility, ensuring exam integrity and fairness in remote settings remains a major challenge. A core difficulty lies in effective exam invigilation. In contrast to traditional in-person exams, where instructors can physically monitor students, online proctoring typically requires supervising dozens of simultaneous video streams displayed as small windows in a gallery view. Manual monitoring under such conditions is inherently limited by human attention, prone to fatigue and bias, and does not scale well for large classes or extended exams. Although commercial automated proctoring systems employ computer vision and AI to address these issues, they often rely on intrusive data collection, centralized cloud processing, or complex installations, raising privacy concerns and limiting deployment in resource-constrained or privacy-sensitive environments [proctoru_official_website, balash2021examining, balash2023educators]. As a result, many institutions continue to reply on manual video monitoring despite its known shortcomings.
To address these limitations, we present iExam, a practical and privacy-preserving online exam proctoring system that combines lightweight real-time face detection with deep learning-based post-exam analysis. The key design insight behind iExam is to decouple instant invigilation from in-depth integrity verification, recognizing that these stages impose fundamentally different requirements. During the exam, real-time proctoring demands low latency, scalability and minimal resource consumption. After the exam, retrospective analysis prioritizes accuracy and richer behavioral assessment. As shown in Fig. 1, iExam adopts a two-phase architecture. During the exam, a lightweight face detection module runs entirely in the browser capturing the Zoom gallery view. This enables immediate feedback, such as alerts for face absence, while preserving privacy by avoiding the transmission of sensitive video data to external servers. After the exam, recorded video streams are processed offline using Deep Convolutional Neural Networks (DCNNs) [krizhevsky2012imagenet] for face recognition, enabling robust detection of sophisticated anomalies such as identify substitution or repeated face disappearance. To eliminate manual annotation, iExam further introduces an enhanced Optical Character Recognition (OCR)-based labeling pipeline that reliably extracts student identities from dynamic on-screen name labels. To ensure practical deployability on standard teacher laptops, iExam incorporates efficiency-oriented optimizations, including sparse OCR labeling and automated model selection, significantly reducing computational overhead without sacrificing accuracy. Extensive experiments on real-world online exam datasets demonstrate that iExam achieves 90.4% accuracy in real-time face detection and 98.4% accuracy in post-exam recognition, while producing visual reports that reduce instructor workload and improve transparency.
Contributions. In sum, our contributions are follows:
@@
  • We design and implement iExam, a two-stage online exam proctoring system that integrates privacy-preserving real-time face detection and deep learning-based post-exam face recognition and anomaly analysis.
@@
  • We propose an automated OCR-based labeling pipeline to extract student identities from dynamic Zoom window labels without manual annotation.
  • We introduce resource-efficient training and inference optimizations that enable effective deployment on standard teacher hardware.
@@
  • We validate iExam on real-world online exam videos, demonstrating high accuracy and actionable reports for scalable and trustworthy exam supervision.
@@
II Lightweight Real-time Video Monitoring
@@
A key advantage of iExam is its privacy-preserving, fully client-side design. All face detection is performed locally within the browser; no video frames or facial features are ever transmitted externally. This edge-based architecture eliminates privacy risks inherent to cloud processing and enables seamless, scalable deployment in real-world educational settings.
User Workflow. During an online exam, invigilators open the web-based iExam platform and select the Zoom window to monitor. The system mirrors the chosen window and overlays real-time face detection results, displaying both the original and processed video streams. The face detection model is loaded on demand, and all computation occurs locally, ensuring fast and seamless operation without external dependencies.
@@
III Comprehensive Post-exam Video Analysis
@@
The workflow of this component consists of three main stages: (1) face dataset construction, (2) classifier training, and (3) face monitoring and feedback generation. To initiate the process, the proctor provides a short video segment (e.g., five minutes) sampled from the recorded exam session, which is used to train individualized face recognition models for that exam. This training is performed once per exam to adapt to session-specific visual conditions before analyzing the full recording.
@@
In this task, two technical challenges must be addressed:
@@
III-A Automatic Ground Truth Collection
@@
The Limitation of Static Position Labeling. The Zoom gallery mode arranges participants in a $5\times 5$ grid, but their order is not fixed. As students join, leave, or lose connection, the positions are shuffled: newcomers are usually placed at the lower right, while others move to fill vacant spots. Thus, it is impossible to rely on a static mapping between cell position and student identity.
@@
III-B Effective and Efficient Face Training and Recognition
@@
Model Training and Selection. We evaluated a range of state-of-the-art DCNN architectures, including AlexNet [krizhevsky2012imagenet], GoogleNet [szegedy2015going], various ResNet variants (ResNet18, ResNet50, ResNet152) [he2016deep], as well as SqueezeNet [SqueezeNet], MobileNet [howard2017mobilenets], and DenseNet121 [huang2017densely], using cross-entropy loss as the training objective. To mitigate overfitting arising from the temporal coherence of consecutive video frames, we performed data augmentation such as random flipping, rotation, and cropping during training. This increases the diversity of training samples and enhances the model’s generalization ability.
@@
IV Evaluation
IV-A The Performance and Accuracy of Face Detection Algorithms in Real-time Exam Monitoring
@@
Experimental Setup. We collected data from an actual online exam session, resulting in a two-hour video featuring the Zoom gallery view. For this experiment, we implemented two versions of iExam’s real-time monitoring system: one using the traditional Haar-cascade face detector, and the other employing a modern DCNN-based detector. Apart from the face detection module, all other components and settings, including browser environment and platform configurations, remained identical.
@@
Runtime Performance. We first measured the detection efficiency of each face detection algorithm, focusing on the number of successful detections and the corresponding processing time. As shown in Fig. 5(a), Haar-cascade achieves about twice the detection rate of DCNN on the same video stream, primarily due to its simpler, feature-based approach and lower computational requirements. However, this higher speed comes at the expense of accuracy and robustness, as the Haar-cascade method is more prone to missing faces under challenging conditions (e.g., large head rotations, partial occlusions). In contrast, while the DCNN-based detector processes frames at a lower rate, requiring roughly twice as long to analyze each image, it consistently provides more reliable detection results, especially in cases where faces are tilted, partially covered, or presented at non-frontal angles. This robustness is attributed to the learning-based nature of DCNNs, which enables them to generalize beyond the fixed feature templates used in traditional methods.
Detection Accuracy. To quantify detection accuracy, we conducted a manual evaluation by randomly sampling 50 frames from the two-hour video and annotating the ground truth for each student’s face. Fig. 5(b) illustrates the results, showing that the DCNN detector achieves around 90% detection accuracy, significantly outperforming the Haar-cascade method. Notably, the DCNN approach is better able to adapt to a variety of challenging viewing conditions, such as faces appearing close to the camera, rotated, or partially obscured. In contrast, Haar-cascade often fails to detect faces that deviate from standard frontal poses or are occluded, and sometimes misidentifies background objects as faces.
@@
IV-B The Performance and Accuracy of Face Recognition Models in Post-exam Video Analysis
@@
Experimental Setup. We first partition the available face image dataset into three disjoint subsets: 70% for training, 20% for validation, and the remaining 10% for testing. All file indices are randomly shuffled prior to splitting to avoid selection bias. To further verify the robustness of our models, we also evaluate on an independent 90-minute exam video not used during training. All training experiments were performed on a laptop equipped with an NVIDIA GTX 1650Ti GPU (4GB VRAM), simulating a practical deployment environment for educational institutions.
Training Performance. We systematically compare the training time and recognition accuracy of eight popular DCNN architectures, including AlexNet, GoogleNet, ResNet18, ResNet50, ResNet152, SqueezeNet, MobileNet, and DenseNet121. As shown in Fig. 6, GoogleNet stands out as the fastest model, requiring only 244 minutes to complete 10 epochs of training on our dataset. AlexNet and ResNet18 also demonstrate relatively fast training speeds, completing in 293 and 303 minutes, respectively. In contrast, deeper architectures such as ResNet152 and SqueezeNet require substantially more time, 650 and 555 minutes, respectively, reflecting the increased computational complexity of larger models. MobileNet and DenseNet121 fall in the mid-to-upper range, with training times of 469 and 445 minutes, respectively.
@@
Recognition Accuracy. Despite the differences in training efficiency, we observe notable variation in the face recognition accuracy achieved by each model, as summarized in the Fig. 6. ResNet50 achieves the highest accuracy at 98.41%, establishing a clear lead over other models. ResNet152 and GoogleNet also perform well, reaching 93.02% and 88.28% accuracy, respectively. AlexNet and MobileNet deliver moderate results, with accuracies of 77.26% and 78.33%. SqueezeNet lags behind with the lowest accuracy at 69.07%, likely due to its lightweight architecture and reduced parameter count. DenseNet121 and ResNet18 achieve intermediate performance (86.17% and 83.82%, respectively).
@@
(a) Overall recognition frequency
(b) Temporal analysis
(c) Conclusion of consecutive absence
@@
Figure 7: Visual reports generated by iExam, showing (a) the number of times each student’s face was detected and identified during the exam, (b) a summary of recognition distribution indicating periods when students were present or absent, and (c) the longest continuous absence interval for each student. Real names have been replaced with generic labels to protect privacy.
IV-C A Case Study of iExam’s Post-exam Analysis Report
To demonstrate the practical utility of iExam’s post-exam analysis, we present a case study using a five-minute exam video. The system automatically generates histogram-style visual reports that allow instructors to quickly review student attendance and detect abnormal behaviors.
Recognition Frequency Overview. As shown in Fig. 7(a), the report includes a bar chart summarizing the total number of times each student’s face was recognized. To improve efficiency, one successful recognition per student per second is counted as a valid presence record, rather than counting every frame. This reduces computation while providing a reliable measure of engagement, helping identify students whose faces were rarely detected and potentially flagging technical issues or suspicious behavior.